OpenStudy (anonymous):

a train, x meters long, traveling at a constant speed, takes 20 seconds from the time it enters a tunnel 300 meters long until the time it completely emerges from the tunnel. one of the stationary ceiling lights in the tunnel is directly above the train for 10 seconds. find x

OpenStudy (anonymous):

oh i got this one its: 300+x=2x x=300

OpenStudy (anonymous):

i think..... well atleast thts wat it said on yahoo answers.... O_o"

OpenStudy (lilg132):

haha

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Let x be the length of the train in km and r be the rate in km/hr. Solve the following for x and r:\[\left\{r \frac{20}{3600}=x+\frac{300}{1000},r \frac{10}{3600}=x\right\} \]\[\left\{r\to 108 \text { km}/\text{hr},x\to \frac{3}{10}\text{km}\right\} \]